OpenStudy (moazzam07):

Shyla used a probability simulator to pull 3 colored marbles from a bag and flip a coin 50 times. The results are shown in the tables below:

OpenStudy (anonymous):

tables?

OpenStudy (moazzam07):

Using Shyla's simulation, what is the probability of pulling a blue marble and the coin landing tails up? 512 over 2500 624 over 2500 32 over 50 48 over 50

OpenStudy (anonymous):

what is the probability of pulling a blue marble according to the table

OpenStudy (moazzam07):

16/50

OpenStudy (anonymous):

correct, now what is the probability of getting a tails up according to the table

OpenStudy (moazzam07):

32/50

OpenStudy (anonymous):

do you know what to do now?

OpenStudy (moazzam07):

yes

OpenStudy (anonymous):

what do you do?

OpenStudy (moazzam07):

Multiply the probabilities

OpenStudy (moazzam07):

Its A

OpenStudy (moazzam07):

512/2500

OpenStudy (anonymous):

yay

OpenStudy (moazzam07):

Yah

OpenStudy (moazzam07):

Next?

OpenStudy (anonymous):

\[\frac{ 32 }{ 50 }\times \frac{ 16 }{ 50}\]

OpenStudy (anonymous):

next question